Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 120     Entry time: 02 Apr 2023, 04:46
Author: Benjamin van Kuiken 
Group: SCS 
Subject: Late Shift Summary 

Achievement

  • Measured tXAS at the Fe L-edge of Fe(bpy)3 using both delay and energy scans. 

Issues

  • The cold traps ran empty on the chamber resulting in a pressure spike that interlocked the DSSC
  • The rough vacuum of the chem DPS degraded causing TP5 to go into error. This resulted in valve closing and stopping the measurements
  • Jet froze twice
ELOG V3.1.4-7c3fd00